OpenStudy (anonymous):

find x so that a is parallel to b. (4x+11) , (8x+1)

OpenStudy (anonymous):

which is a which is b?

OpenStudy (anonymous):

8x+1

OpenStudy (anonymous):

so you're looking for an x so (4x+11) = 8x+1?

OpenStudy (anonymous):

yes the varies answers are 42, 14, 7

OpenStudy (anonymous):

so 4x - 8x, then 11 + 1, so -4x = 12, then x = -3, is it right?

OpenStudy (anonymous):

no that's not right

OpenStudy (anonymous):

i am confused

OpenStudy (anonymous):

what class is this?

OpenStudy (anonymous):

geometry

OpenStudy (anonymous):

its a virtual school

OpenStudy (anonymous):

have you graphed these?

OpenStudy (anonymous):

it is three lines crossing one long line and a crossing the long line at one point and b crossing at another a and b never touch

OpenStudy (anonymous):

I'm thinking one line is y=4x+11, and another is y=8x+1. Is this true? If so, which line is A?

OpenStudy (anonymous):

4x+11

OpenStudy (anonymous):

if that is the case then how can these lines ever be parallel...........????

OpenStudy (anonymous):

i dont know